OpenStudy (anonymous):

A child is creating a pyramid with building blocks. The top three levels include 5 blocks, 12 blocks, and 19 blocks. How many blocks would be needed for a pyramid 25 levels tall? Use complete sentences to explain how a sum of an arithmetic series was applied?

OpenStudy (anonymous):

i would use the S=n/2[2a+(n-1)d] formula but im having problems with wat n would be ? a is 5 right?

OpenStudy (anonymous):

part1, a25 = 5+(24)7=173, let a1=5,a25=173 then solve for the sum s = (25(5+173))/2 = 2225,answer part2,let 5 be our 1st term, 173 be our last term and 25 be our number of terms then used the formula

OpenStudy (anonymous):

wat formula is that ? is that a + (n-1 )d ?